OpenStudy (anonymous):

ind the eqn of tangent line to x^2 + 2xy - y^2 + x = 2 at (1,2) Im not sure how you would isolate y and take the derivative

OpenStudy (anonymous):

this is a quadreatic in y solve using the quadratic equation. i'll try to get an answer to you. computer is very slow

OpenStudy (anonymous):

first solve the initial quadratic in y , (-y^2 +2xy-(2-x+x^2)=0) which should give you -x +- root(2x^2-x+2) =y you should be able to differentiate it from here knowing that a root is the same as to the power of a half.

OpenStudy (anonymous):

\[2x+2y+2x \frac{dy}{dx}-2y \frac{dy}{dx}+1=0\] \[2x+2y+1=\frac{dy}{dx} (-2x+2y)\] \[\frac{dy}{dx}=\frac{2x+2y+1}{-2x+2y}\]

OpenStudy (anonymous):

Now plug in (1,2) \[\frac{2(1)+2(2)+1}{-2(1)+2(2)}=\frac{7}{2}\]

OpenStudy (anonymous):

If I remember implicit differentiation correctly haha.

OpenStudy (anonymous):

@jf55 You don't have to isolate the y to differentiate it. You can do whats called implicit. Think about this, if you have y=f(x) And you differentiate it with respect to x, you get: \[\frac{dy}{dx}=f'(x)\] Its the same thing if you have a y in the function. \[\frac{d}{dx}(x^2y+2x+y^3)=2xy+x^2 \frac{dy}{dx}+2+3y^2 \frac{dy}{dx}\] (NOTE: the first one you have to use a product rule)

OpenStudy (anonymous):

@jf55 You don't have to isolate the y to differentiate it. You can do whats called implicit. Think about this, if you have y=f(x) And you differentiate it with respect to x, you get: \[\frac{dy}{dx}=f'(x)\] Its the same thing if you have a y in the function. \[\frac{d}{dx}(x^2y+2x+y^3)=2xy+x^2 \frac{dy}{dx}+2+3y^2 \frac{dy}{dx}\] (NOTE: the first one you have to use a product rule)